# Graph Neural Networks for Next-Generation-IoT: Recent Advances and Open Challenges

### v1

@@ III Applications of GNN for NG Technologies in IoT Networks
In this section, we address Open Questions 3 and 4, focusing on how GNNs enhance key NG technologies. Open Question 3 covers resource management and communication efficiency in massive MIMO, RIS, Satellite, MEC, THz, and URLLC systems, with GNNs optimizing user association, beamforming, and signal propagation. By contrast, open Question 4 investigates how GNNs enhance integrity and security in blockchain.
TABLE VI: Related studies in the use of GNN on different NG technologies for IoT networks
@@
Open Question 3: How Do GNNs Enhance Communication and Computation Efficiency in Massive MIMO, RIS, Satellite, THz, MEC, and URLLC Systems?
@@ III-3 Challenges and future directions
Open Question 4: How Can GNNs Enhance Integrity, Security, and Scalability Along with Blockchain for IoT Systems?
@@ IV Adversarial Attacks and Defense Mechanisms Conceived for GNN-Based NG-IoT Networks
In this section, we address a pair of critical questions: Open Question 5 focuses on the nature of adversarial attacks targeting GNN-based systems, while Open Question 6 explores defense techniques designed for safeguarding these systems. We commence by examining how adversarial attacks exploit vulnerabilities in GNN models within NG-IoT networks, targeting both homogeneous GNNs (HoGNNs) and heterogeneous GNNs (HeGNNs). These attacks can severely impact the performance and security of systems in applications such as smart cities, autonomous transportation, and healthcare. Next, Open Question 6 explores defense mechanisms that have been developed for countering these attacks, enhancing both the robustness and reliability of GNN-based systems. Finally, we review the latest strategies of adversarial defense and provide insights into future research directions for securing GNNs in the complex dynamic environments of NG-IoT networks.
Open Question 5: How Do Adversarial Attacks Exploit Vulnerabilities in GNN-Based NG-IoT Networks?
@@ IV-3 Challenges and future directions:
Open Question 6: What Are the Most Effective Defense Techniques Against Adversarial Attacks in GNN-Based NG-IoT Networks?
@@ V The Role of GNNs in Future Integrated Networks and Quantum Computing
Open Question 7: How Can GNNs Enhance the Performance and Scalability of Future Integrated Sensing and Communication and Space-Air-Ground-Sea Integrated Networks?
@@ V-3 Challenges and future directions:
Open Question 8: How Can GNNs and Future Computational Technologies, Like Quantum Computing, Work Together to Enhance the Capabilities of NG-IoT Networks?
@@ VI Generic Design Guidelines
In order to design an efficient GNN model for next-generation IoT networks, the following design guidelines may be considered:
  • Task specific design: GNN models are highly effective for node and edge-level tasks, such as user association, resource allocation, and network routing, providing precise predictions and improved network performance. For larger-scale problems, graph-level tasks offer a holistic view of the network, capable of optimizing, for example, throughput and task offloading in complex environments like edge computing.
@@
  • Balancing performance and computational complexity: The objective is to design lightweight GNNs imposing a computational overhead, while maintaining high performance. Incorporating quantum computing has the potential of enhancing efficiency by offloading computationally intensive tasks to quantum circuits, in large-scale graphs. Besides, the depth of the GNN model must be carefully designed to strike a balance between achieving node feature uniqueness and avoiding over-smoothing.
@@
VII Conclusion
@@
A critical appraisal of the application of Graph Neural Networks (GNNs) in NG-IoT networks was presented. We commenced by introducing the family of GNN paradigms and clarifying the roles and advantages of node, edge, and graph-level tasks in addressing wireless networking challenges. Through practical use cases and examples, we demonstrated how each task level can be effectively applied for improving problem-solving efficiency across various NG-IoT scenarios. We also examined the application of GNNs to key NG technologies, including massive MIMO schemes, RISs, satellites, THz, MEC, and URLLC solutions, highlighting their capability to circumvent the limitations of traditional methods. Furthermore, we explored the integration of GNNs with future technologies, including integrated sensing and communication, space-air-ground-sea integrated networks, and quantum computing, showcasing their potential to enhance NG-IoT networks. Additionally, we addressed a range of critical security concerns by discussing adversarial attacks on GNN-based systems and by appraising the family of effective defense strategies. In conclusion, our findings have emphasized the transformative potential of GNNs in optimizing, scaling, and securing NG-IoT networks, laying the groundwork for future research in this rapidly evolving field.

### v2

@@ III Applications of GNN for NG Technologies in IoT Networks
In this section, we address Open Questions 4, 5, and 6, focusing on how GNNs empower core NG technologies Open Question 4 covers resource management and communication efficiency in massive MIMO, RIS, Satellite, MEC, THz, and URLLC systems, with GNNs optimizing user association, beamforming, and signal propagation. By contrast, Open Question 5 investigates how GNNs enhance integrity and security in blockchain. Lastly, Open Question 6 discusses the integration of GNNs with distributed systems, examining their collaboration to enable privacy-preserving, scalable, and efficient learning across distributed NG-IoT networks.
TABLE VII: Related studies in the use of GNN on different NG technologies for IoT networks
@@
Distributes systems [96] (2024) Federated learning for network attack detection Attack detection Node GAT
[97] (2023) Federated learning Anomaly detection Node GNN
[98] 2024 Federated learning for UAV assisted MEC systems Classification inference Node GNN
[99] (2025) Federated learning for classification tasks Node classification tasks Node GNN
[100] (2025) Federated learning for edge computing Node classification tasks Node RL + GNN
[102] 2025 Decentralized tasks in wireless networks Channel impairments in GNNs Node GNN
[103] 2022 Decentralized federated learning in V2V system Power control Node GCN
[104] 2024 Distributed fog servers for smart IoT services Communication efficiency Node GNN
[105] 2024 Federated learning over interference-limited wireless networks Power control Node GCNs
[38] (2023) RIS support federated learning Resource allocation Node GNN
[106] (2023) Federated learning over wireless network Power control Node GNN
[92] (2024) Distributed URLLC system Beamforming and power control Node GNN
Open Question 4: How Do GNNs Enhance Communication and Computation Efficiency in Massive MIMO, RIS, Satellite, THz, MEC, and URLLC Systems?
@@ III-3 Challenges and future directions
Open Question 5: How Can GNNs Enhance Integrity, Security, and Scalability Along with Blockchain for IoT Systems?
@@ III-3 Challenges and future directions:
Open Question 6: How Can GNNs Be Integrated with Distributed Systems in NG-IoT Networks?
III-1 Backgrounds:
As NG-IoT environments consist of diverse, distributed, and often resource-constrained edge devices, centralized data collection and processing pose critical challenges in terms of data privacy, communication costs, and model scalability. To address these issues, distributed paradigms, such as federated learning (FL) and decentralized structures, have gained significant attention. FL allows multiple edge devices to collaboratively train a shared model without exchanging raw data, preserving privacy and reducing communication overhead [182, 183, 184, 185, 186, 187]. Meanwhile, decentralized architectures eliminate the need for a central server by enabling peer-to-peer collaboration among devices [188, 63]. The integration of GNNs with distributed systems has emerged as a promising paradigm to meet the privacy, scalability, and efficiency requirements of next-generation IoT (NG-IoT) networks. Combining these two paradigms empowers edge nodes in NG-IoT systems to collaboratively learn graph-based models with minimal communication overhead and stronger privacy guarantees. This synergy opens two prominent research directions: (1) leveraging distributed architectures to optimize GNN(a) and (2) using GNNs to enhance distributed system in NG-IoT systems. Fig. 12 illustrates the integration of GNN with distributed frameworks in NG-IoT: (a) Federated learning-assisted GNN, where a central server aggregates local GNN models trained on distributed clients; (b) GNN-enabled decentralized system, where local agents make decisions collaboratively based on peer-to-peer graph reasoning without a central server.
III-2 State-of-the-art:
Figure 12: The incorporation of GNN and FL: (a) FL-assisted GNN, and (b) GNN-assisted a decentralized system
Distributed architecture for GNN in NG-IoT networks: Distributed systems, especially FL, provide robust frameworks for collaboratively training GNNs across distributed nodes while preserving privacy. For instance, Wu et al. [96] developed FedGAT, a federated graph attention network for privacy-preserving network attack detection, leveraging attention to improve internal traffic interaction modeling. Similarly, Zhang et al. [97] employed FL to train a GNN-based anomaly detection system for controller area networks (CAN bus), preserving privacy without compromising detection accuracy. Addressing resource constraints in mobile edge computing (MEC), Zhong et al. [98] proposed a lightweight federated GNN framework for UAV-assisted environments, combining adaptive information bottlenecks and compact GNNs for accelerated classification. Meanwhile, Huang et al. [99] optimized both client and cross-client edge selection to improve convergence performance and cost-efficiency in training GCNs over distributed data. Khanna et al. [100] combine FL with graph reinforcement learning to optimize edge caching in MEC, demonstrating latency and cache hit improvements.
In [101], Li et al. presented DHTGL, a distributed hierarchical temporal graph learning framework, reducing communication in IIoT systems. For decentralized training environments, Gao et al. [102] introduced AirGNNs, a communication-aware GNN architecture that incorporates wireless channel effects and enables over-the-air decentralized GNN training. The framework is robust to fading and noise and supports decentralized tasks such as multi-robot flocking and source localization. Li et al. [103] proposed a GCN-based dynamic topology selection strategy in vehicle-to-vehicle (V2V) networks to ensure reliable model sharing in decentralized federated learning (DFL), thereby improving training stability in mobile environments. In [104], Zeng et al. proposed FoGraph, a distributed GNN inference platform deployed over fog nodes. By leveraging compression-aware execution planning, FoGraph reduces communication bottlenecks compared to traditional cloud-based model serving and accelerates GNN inference by over 5×5$\times$5 ×.
GNN for the distributed systems in NG-IoT networks: Several studies utilize GNNs to improve the design and performance of decentralized systems and FL. Li et al. [105] proposed a GCN-based power allocation policy designed for wireless FL systems operating under channel variability and non-IID data distributions. Their method parameterizes power control using a GCN and solves the associated optimization via a primal-dual approach, ensuring communication efficiency while maintaining learning accuracy. Wang et al. [38] incorporated GNNs into over-the-air FL by learning efficient mappings from wireless channel states to RIS-assisted transceiver configurations, thus mitigating aggregation bottlenecks and improving spectral efficiency. In [106], Yang et al. introduced a two-stage GNN-based collaborative FL framework aimed at minimizing energy consumption. The model enables each device to learn an optimal set of communication links for local parameter exchange and corresponding transmit power levels, significantly outperforming traditional iterative optimization techniques in both speed and energy efficiency. In another work, Zhang et al. [92] leveraged GNN-based policies to solve the joint distributed beamforming and power control problem in massive URLLC networks, targeting a reduction in decoding error probabilities under tight latency constraints.
III-3 Challenges and future directions:
Despite these advances, several challenges remain in integrating GNNs with the distributed systems in IoT networks. First, heterogeneity in data and system conditions, such as device capabilities, communication channels, and data distributions, restricts effective collaboration. Besides, devices in NG-IoT systems possess different capabilities in computational power, memory, and energy constraints. This results in non-uniform convergence behavior and degraded model performance in federated GNN training, especially in highly dynamic environments like UAVs, mobile edge networks, or vehicular systems.
Another significant problem is about communication efficiency, although approaches such as over-the-air computation (e.g., AirGNNs [102], Air-MPNN [189]) and graph compression techniques (e.g., LW-FGL [98], DHTGL [101]) aim to mitigate communication overhead, the scalability of these methods in extremely large or densely connected networks is still underexplored. Besides, distributed nodes requires synchronization and consistency, which remains difficult to achieve when topology evolves rapidly (e.g., in V2V networks), requiring more robust protocols for model aggregation and graph updates. Finally, dynamic graph learning under limited feedback and partial observability remains a key challenge, as NG-IoT graph structures often evolve due to mobility and device failures. Future GNN architectures must support online updates and continual learning without full graph re-computation to meet real-time demands.
@@ IV Adversarial Attacks and Defense Mechanisms Conceived for GNN-Based NG-IoT Networks
In this section, we address a pair of critical questions: Open Question $9$ focuses on the nature of adversarial attacks targeting GNN-based systems, while Open Question $10$ explores defense techniques designed for safeguarding these systems. We commence by examining how adversarial attacks exploit vulnerabilities in GNN models within NG-IoT networks, targeting both homogeneous GNNs (HoGNNs) and heterogeneous GNNs (HeGNNs). These attacks can severely impact the performance and security of systems in applications such as smart cities, autonomous transportation, and healthcare. Next, Open Question 6 explores defense mechanisms that have been developed for countering these attacks, enhancing both the robustness and reliability of GNN-based systems. Finally, we review the latest strategies of adversarial defense and provide insights into future research directions for securing GNNs in the complex dynamic environments of NG-IoT networks.
Open Question 7: How Do Adversarial Attacks Exploit Vulnerabilities in GNN-Based NG-IoT Networks?
@@ IV-3 Challenges and future directions:
Open Question 8: What Are the Most Effective Defense Techniques Against Adversarial Attacks in GNN-Based NG-IoT Networks?
@@ V The Role of GNNs in Future Integrated Networks and Quantum Computing
Open Question 9: How Can GNNs Enhance the Performance and Scalability of Future Integrated Sensing and Communication and Space-Air-Ground-Sea Integrated Networks?
@@ V-3 Challenges and future directions:
Open Question 10: How Can GNNs and Future Computational Technologies, Like Quantum Computing, Work Together to Enhance the Capabilities of NG-IoT Networks?
@@ VI Generic Design Guidelines
Figure 18: Design guidelines for efficient GNNs in NG-IoT networks.
To guide the development of efficient GNN models tailored for NG IoT networks, we highlight key design principles, as summarized in Fig. 18, and elaborated below.
  • Task-specific design: GNN models are highly effective for node and edge-level tasks, such as user association, resource allocation, and network routing, providing precise predictions and improved network performance. For larger-scale problems, graph-level tasks offer a holistic view of the network, capable of holistic system optimization, for example, throughput and task offloading in complex environments like edge computing.
    Evaluation metrics: The effectiveness of this design can be validated using task-specific performance metrics, e.g., accuracy, precision, or latency for edge-level tasks; and throughput, fairness index, or load balance for graph-level tasks. These metrics indicate whether the GNN architecture is aligned with the intended scope of decision-making, ensuring that the model does not underfit or overfit the level of abstraction.
@@
    Evaluation metrics: Metrics such as graph entropy, equivariance, invariance, and attention weight distribution provide valuable insights into how effectively a GNN model captures and preserves the underlying structural information. Additionally, inference accuracy under partial graph observation evaluates the model’s ability to propagate critical information across the network, while EXIT chart-inspired metrics can monitor mutual information flow between layers, helping diagnose issues such as over-smoothing or representation collapse. Overall, these metrics demonstrate the critical role of representational power in achieving robust performance in complex NG-IoT environments.
@@
    Evaluation metrics: The key metric is scalability-efficiency, measuring how performance (e.g., accuracy or reward) changes with varying graph sizes. This indicator is essential for validating the model’s practicality in real-time applications and resource-constrained NG-IoT deployments, ensuring it remains effective as the network scales.
  • Balancing performance and computational complexity: The objective is to design lightweight GNNs imposing a low computational overhead, while maintaining high performance. Incorporating quantum computing has the potential of enhancing efficiency by offloading computationally intensive tasks to quantum circuits, in large-scale graphs. Besides, the depth of the GNN model must be carefully designed to strike a balance between achieving node feature uniqueness and avoiding over-smoothing.
    Evaluation metrics: Trade-off metrics like accuracy-per-FLOP, model size vs. latency, and energy-per-inference provide quantitative insights into this balance. These metrics will evaluate the deployment feasibility on constrained hardware (e.g., FPGAs or Edge TPUs), hence facilitating compact and effective GNN designs in NG-IoT networks.
@@
    Evaluation metrics: The main evaluation metrics will be accuracy drop (DA%), perturbation tolerance thresholds, and resistance to evasion or poisoning attacks. Evaluating model reliability under simulated threats provides empirical justification for integrating defense mechanisms, proving that the model can maintain trustworthiness in volatile or adversarial environments.
VII Lessons Learned, Future Research Direction, and Conclusions
VII-A Lessons Learned and Future Research Directions
Based on the comprehensive discussion of each open question, we summarize the key lessons learned and highlight promising future research directions to further strengthen the integration of GNNs with NG-IoT networks.
  • GNNs can provide a flexible and expressive modeling tool to capture spatial, temporal, and structural dependencies in NG-IoT systems, enabling intelligent decision-making across edge, cloud, and hybrid platforms. However, the integration of GNNs into these systems is challenged by scalability limitations, heterogeneous data, and the constrained resources of IoT devices. Therefore, future researches should focus on developing lightweight, adaptive GNN architectures, using techniques such as model pruning, quantization, and dynamic graph sparsification, enabling real-time and energy-efficient deployment in practical NG-IoT scenarios.
  • The involvement of multimodal and non-Euclidean data with complex interdependencies causes traditional ML models ineffective in capturing topological structures and dynamic interactions. Meanwhile, GNNs are naturally suited to learn from these structures, but their application to real-time and distributed settings still requires innovation in data fusion strategies, synchronization mechanisms, and dynamic graph representation learning. This requires a robust multimodal GNN pipelines with attention-based fusion, continual learning capabilities, and temporal adaptation for evolving graphs.
  • From a system-level perspective, the deployment of GNNs at the edge introduces a trade-off between performance, energy, and cost. We learn that embedding GNNs into edge devices (e.g., FPGAs, TPUs) and optimizing them through compression and hardware-software co-design can dramatically reduce latency and communication overhead. This necessitates research into green AI strategies, edge-aware GNN design, and scalable training techniques for large-scale IoT environments. Moreover, real-world adoption also depends on developing standardized frameworks for benchmarking GNN models in terms of accuracy, efficiency, and sustainability across heterogeneous NG-IoT settings.
  • Distributed systems offer a natural fit for GNNs in NG-IoT environments where data privacy, communication cost, and scalability are critical. Yet, federated GNN training faces challenges from device heterogeneity, partial observability, and synchronization issues in dynamic graphs. This highlights the need for new protocols and architectures supporting structure-aware aggregation, over-the-air computation, and robust federated updates in non-IID, time-varying settings. Besides, an unified frameworks and public benchmarks will be essential to accelerate reproducibility and progress in this domain.
  • Security and robustness remain pressing concerns. GNNs are vulnerable to adversarial attacks that manipulate graph structures or features, threatening reliability in mission-critical NG-IoT applications. Existing defense techniques provide partial solutions, but many are tailored to specific attack models or fail under dynamic graph conditions. Future directions should include designing universally robust GNNs using adversarial training, explainable AI, and quantum-safe encryption techniques, especially for sensitive applications in smart grids, autonomous transport, and blockchain-integrated systems.
  • In emerging applications like ISAC and SAGSINs, GNNs show promise in managing massive interconnectivity and optimizing cross-domain coordination. However, challenges in handling highly dynamic topologies, spatial-temporal dependencies, and extreme heterogeneity remain. Future work should investigate hierarchical and modular GNN models that can seamlessly operate across different communication layers and adapt to time-varying resources.
  • Finally, quantum computing opens new frontiers in accelerating GNN training, enhancing privacy via quantum cryptography, and expanding computational capacity for large-scale NG-IoT graphs. Despite limitations of current noisy quantum hardware, hybrid quantum-classical GNN models and quantum-enhanced optimization can offer long-term benefits. Additionally, designing scalable QGNN frameworks is critical to fully harness the potential of quantum computing for large-scale NG-IoT applications.
VII-B Conclusions
@@
A critical appraisal of the application of Graph Neural Networks (GNNs) in NG-IoT networks was presented. We commenced by introducing the underlying fundamental graph concepts and GNN paradigms, establishing the foundational principles necessary for understanding their relevance. Building on this, we highlighted the pivotal role of GNNs in NG-IoT networks and modeled them as dynamic constrainted graphs. To further justify the GNN’s applicability, we related them to LDPC decoding, framing GNNs as learnable generalizations of belief propagation, hence supporting convergence and information flow analysis via density evolution and EXIT charts. We further discussed how selecting appropriate task levels, among node-, edge-, or graph-level, can significantly enhance problem-solving efficiency across diverse NG-IoT scenarios. We also examined the application of GNNs to key NG technologies, including massive MIMO schemes, RISs, satellites, THz, MEC, and URLLC solutions, highlighting their capability to circumvent the limitations of traditional methods. We also analyzed their integration with distributed systems, highlighting their role in supporting distributed intelligence with privacy preservation and reduced communication overhead. Furthermore, we explored the integration of GNNs with emerging technologies, including integrated sensing and communication, space-air-ground-sea integrated networks, and quantum computing, showcasing their potential to enhance NG-IoT networks. Additionally, we addressed a range of critical security concerns by discussing adversarial attacks on GNN-based systems and by appraising the family of effective defense strategies. By systematically addressing a suite of ten open questions, this survey fills a crucial gap in understanding how GNNs can be efficiently leveraged in NG-IoT networks. In conclusion, our findings have emphasized the transformative potential of GNNs in optimizing, scaling, and securing NG-IoT networks, laying the groundwork for future research in this rapidly evolving field.